Find the Maclaurin series or Taylor series centered at a, interval of

convergence and the radius of convergence of 1

√1−𝑥

.


1
Expert's answer
2021-08-24T15:58:48-0400

Solution;

Maclaurin series is given by;

Since a=0;

"f(x)=\displaystyle\sum_{n=0}^\infin\frac{f^n(0)x^n}{n!}=f(0)+f'(0)x+f"(0)\\frac{x^2}{2!}+f'''(0)\\frac{x^3}{3!}+..."

Given;

"f(x)=\sqrt{1-x}"

"f(0)=\sqrt{1-0}=1"

By differentiation;

"f'(x)=-\frac{1}{2\sqrt{1-x}}"

"f'(0)=-\frac{1}{2\sqrt{1-0}}=-\frac12"

"f''(x)=-\frac{1}{4(1-x)^{\frac32}}"

"f''(0)=-\frac{1}{4(1-0)^{\frac32}}=-\frac14"

"f'''(x)=-\frac{3}{8(1-x)^{\frac52}}"

"f'''(0)=-\frac{3}{8(1-0)^{\frac52}}=-\frac38"

By substitution;

"f(x)=1-\frac12x-\frac18x^2-\frac1{16}x^3+..."

From the series;

"f(x)=\displaystyle\sum_{n=1}^{\infin}(-1)\frac{x^n}{2(n^{2n-2})}"

"a_n=-\frac{1}{2(n^{2n-2}}"

"a_{n+1}=\frac{-1}{2(n+1)^{2n}}"

Using root test;

"L=|x|\displaystyle{lim_{n\to\infin}}|\frac{a_{n+1}}{a_n}|"

"L=|x|lim_{n\to\infin}|\frac{n^{2n-2}}{(n+1)^{2n}}|"

"L=|x|"

The interval of convergence will be ;

"1<x<1"

Hence Radius of convergence is;

"R=1"